Bob, single, has total income of $120,000 in 2020. His taxable income is $100,000, and his tax liability is $18,079.50. Bob's average tax rate is . , and his effective tax rate is . . Bob's marginal tax rate is . . Note: Enter your answer as a decimal rounded up to three places (e.g. 23.175231% would be entered as .232 and 39% would be entered as .390). Do not enter the "%" sign.
Answer #1

Bob’s average tax rate = total tax / total taxable income = 18079.50/100000 = 0.181

Bob’s effective tax rate = total tax / total income = 18079.50/120000 = 0.151

Bob’s marginal tax rate = 0.240

(According to 2020 tax rates, $100000 falls in the tax bracket of 24% for single tax return filer)